Directions: Evaluate the ff functions.



1.p(a) =4³a find p(1)



2.a(c) = 2c + c²-2 find c(-1)



3.f(x)=x² -3x find f (-8)



4.b(d) =-4d² find b(-13)



5.h(n) =-2h²+4 find h(4)

1
Expert's answer
2022-03-07T23:42:06-0500

1) "p(1)=4^3*1=64."

2) "a(-1)=2(-1)+(-1)^2=-1."

3) "f(-8)=(-8)^2-3(-8)=88."

4) "b(-13)=-4(-13)^2=-676."

5) "h(4)=-2*4^2+4=-28."